OpenStudy (anonymous):

7th grade math help PLeaSe

OpenStudy (anonymous):

Which literal equations are equivalent to k= F/x? Choose all answers that are correct. A. F= kx B. x= kf C. F= k/x D. X= F/k

ganeshie8 (ganeshie8):

k= F/x cross multiplying you get : ` kx = F `

OpenStudy (anonymous):

U can choose more then one:)

ganeshie8 (ganeshie8):

compare your options against this just cross multiply for options with fractions

ganeshie8 (ganeshie8):

Oh yah go one by one

OpenStudy (anonymous):

Which literal equation is equivalent to 1.4 = kx